# Causal Effect Estimation under Networked Interference without Networked Unconfoundedness Assumption

### v1

@@ 4 Networked Causal Effect Identification via Representation Learning
To begin with, following existing identifiable representation learning methods (Khemakhem et al., 2020; Lu et al., 2022), we first introduce the generative model as follows:
$p_{\bm{\theta}}(X,U|X_{\mathcal{N}})$ $=p_{\bm{f}}(X|U)p_{\bm{T},\bm{\lambda}}(U|X_{\mathcal{N}})$ (2)
$p_{\bm{f}}(X|U)$ $=p_{\bm{\epsilon}}(X-\bm{f}(U))$
@@
Further, we assume $p_{\bm{T},\bm{\lambda}}(U|X_{\mathcal{N}})$ follows the exponential family distribution.
Assumption 4.1.
@@
where $\mathcal{Q}$ is the base measure, $\mathcal{C}$ is the normalizing constant. The $\bm{\lambda}(X_{\mathcal{N}})$ is an arbitrary function, and the sufficient statistics $\bm{T}(U)=[\bm{T}_{f}(U)^{T},\bm{T}_{MLP}(U)^{T}]^{T}$ contains a) the sufficient statistics $\bm{T}_{f}(U)^{T}=[\bm{T}_{1}(U_{(1)})^{T},\dots,\bm{T}_{1}(U_{(d_{U})})^{T}]$ of a factorized exponential family, where all the $\bm{T}_{i}(U_{(i)})$ have dimension larger or equal to $2$ and $d_{U}$ is the dimension of $U$, and b) the output $\bm{T}_{MLP}(U)$ of a neural network with ReLU activations.
This assumption is introduced by Lu et al. (2022). The distribution in Assumption 4.1 is more flexible than the standard assumed distribution condition in identifiable representation learning (Eq. (7) in Khemakhem et al. (2020)). This assumption allows for the case that the different elements of latent confounders are not independent given the conditional set. The term $\bm{T}_{MLP}(U)$ does capture arbitrary dependencies between latent variables since the neural network with ReLU activation has universal approximation power.
@@
Suppose Assumption 4.1 holds, and suppose the following conditions hold: (1) The set $\{X\in\mathcal{O}|\varphi_{\bm{\epsilon}}(X)=0\}$ has measure zero where $\varphi_{\bm{\epsilon}}$ is the characteristic function of density $p_{\bm{\epsilon}}$. (2) $\bm{f}$ is injective and has all second-order cross derivatives. (3) The sufficient statistics in $\bm{T}_{\bm{f}}$ are all twice differentiable. (4) There exist $k+1$ distinct values $x_{\mathcal{N}_{0}},...,x_{\mathcal{N}_{k+1}}$ such that the matrix
$L=($ $\bm{\lambda}(x_{\mathcal{N}_{1}})-\bm{\lambda}(x_{\mathcal{N}_{0}}),$
$\bm{\lambda}(x_{\mathcal{N}_{k+1}})-\bm{\lambda}(x_{\mathcal{N}_{0}}))$
@@
of size $k\times k$ is invertible where $k=|U^{i}|+|U^{c}|+|U^{n}|$ is the dimension of latent variables. Then we learn the true latent variable $U^{i},U^{c},U^{n}$ up to a permutation and simple transformations.
@@
Discussion on assumptions and conditions. Assumption 4.1 indicates our theory holds for a rich family of conditional densities (Wainwright & Jordan, 2008). The assumption on the exponential family distribution is not strong, since many well-known distributions belong to this family, including Gaussian, Uniform, Poisson distributions, and so on. Condition (1)-(4) is a common assumption in representation learning in causal representation learning, e.g., (Khemakhem et al., 2020; Lu et al., 2022). Notably, the most important condition is the condition (4) which requires that the auxiliary information should be sufficient enough. Under our networked setting, it requires that there exist enough distinct values of neighbors’ covariates, which is easy to hold if we collect enough covariates, especially when some of the covariates are continuous.
@@
Theorem 4.2 indicates that, under mild assumptions, the latent confounders can be recovered up to a simple function, i.e., the recovered $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$ satisfying $\hat{U}^{i}=h_{i}(U^{i}),\hat{U}^{c}=h_{c}(U^{c}),\hat{U}^{n}=h_{n}(U^{n})$ for some simple functions $h_{i}.h_{c},h_{n}$. Based on Theorem 4.2 above, we can further identify networked effect as follows:
Theorem 4.3.
Suppose Assumption 3.7, 3.9, 3.8, 3.11, and Theorem 4.2 holds, the networked effects $\psi(t,z)$ and $\mu(t,z,x,x_{\mathcal{N}})$ are identifiable.
@@
Theorem 4.3 indicates that if we can identify latent confounders, the networked effect is thereby identifiable. This necessitates the utilization of identifiable representation learning techniques for causal inference in the presence of networked interference and latent confounders.
@@
Figure 3: Model architecture of our proposed method. The representation learning module aims to learn the unobserved confounders. The feature module aggregates the information of covariates of unit $i$ and its neighbor. The outcome estimator module aims to estimate potential outcomes of unit $i$.
5 Methodology
@@
In this section, leveraging the theoretical findings, we devise our networked effect estimator in the presence of latent confounders. Specifically, our estimator contains three modules, including the representation learning module, the feature module, and the outcome estimator module. The representation learning module is built on Theorem 4.2, aiming to correctly recover three types of latent confounders $u^{i},u^{c},u^{n}$. The feature module is built on Theorem 4.3, aggregating the information from units’ and neighbors’ information. This aggregated information is then input into the outcome estimator module to predict the networked causal effects. Overall, our model architecture is shown in Figure 3.
5.1 Representation Learning Module
@@
Following existing work (Guo et al., 2020; Ma & Tresp, 2021; Jiang & Sun, 2022; Chen et al., 2024), we use Graph Convolution Networks (GCN (Defferrard et al., 2016; Kipf & Welling, 2016)) to aggregate the information of covariates of unit $i$ and its neighbors, i.e., $x_{i},x_{\mathcal{N}_{i}}$:
$h^{neigh}_{i,1}=\sigma(\sum_{j\in\mathcal{N}_{i}}\frac{1}{\sqrt{d_{i}d_{j}}}W_{1}^{T}x_{j}),$
@@
where $\sigma(\cdot)$ is a non-linear activation function, $d_{i}$ is the degrees of unit $i$, $W_{1}$ is the learning weight matrix of GCN, and $MLP_{1}$ is a multilayer perception (MLP).
Then, given $x_{i}$ and $x_{\mathcal{N}_{i}}$, we employ the identifiable representation learning technique (Lu et al., 2022) to recover latent confounders. Specifically, we parametrize the prior following Assumption 4.1:
@@
where $u_{i}=[u_{i}^{i},u_{i}^{c},u_{i}^{n}]$ and $MLP_{2}(u_{i})$ serves as $\bm{T}_{MLP}$, the concatenated $[u_{i},u_{i}^{2}]$ serves as $\bm{T}_{f}$, $MLP_{3}(h_{i,2})$ serves as $\bm{\lambda}_{MLP}$, and $MLP_{4}(h_{i,2})$ serves as $\bm{\lambda}_{f}$ in Assumption 4.1.
@@
$q(u^{i}_{i},u^{c}_{i},u^{n}_{i}|x_{i},x_{\mathcal{N}_{i}})=\prod_{i=0}^{|U|}\mathcal{N}(\mu=\hat{\mu}_{u_{i}},\sigma^{2}=\hat{\sigma}_{u_{i}}^{2}),$ (5)
where $\hat{\mu}_{u_{i}}$ and $\hat{\sigma}_{u_{i}}^{2}$ are the mean and variance of the Gaussian distribution parametrized by MLPs using $h_{i,2}$ as input.
@@
$p(x_{i}|u^{i}_{i},u^{c}_{i},u^{n}_{i})=\prod_{i=0}^{|X|}\mathcal{N}(\mu=\hat{\mu}_{x},\sigma^{2}=v_{x}^{2})$ (6)
or $p(x_{i}|u^{i}_{i},u^{c}_{i},u^{n}_{i})=\prod_{i=0}^{|X|}\mathbf{Bern}(\pi=\hat{\pi}_{x}),$
where for the continuous case $\hat{\mu}_{x}$ is the mean of the Gaussian distribution parametrized by an MLP using the sampled $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ from posterior as input, and $v_{x}^{2}$ is the fixed variance of Gaussian distribution, and for the discrete case $\hat{\pi}_{x}$ is the mean of Bernoulli distribution similarly parametrized by an MLP.
For this module, we use the negative variational Evidence Lower BOund (ELBO) as the loss function, defined as
@@
Following Lu et al. (2022), we utilize the score matching technique (Vincent, 2011) for training unnormalized probabilistic models to learn the parameters of $\bm{T}$ and $\bm{\lambda}$ by minimizing
@@
$\|\nabla_{u}\log q(u^{i},u^{c},u^{n}|x.x_{\mathcal{N}})-\nabla_{u}\log p(u^{i},u^{c},u^{n}|x_{\mathcal{N}})\|^{2}],$
where $\nabla$ is the gradient operator.
5.2 Feature Module and Outcome Estimator Module
@@
where $\sigma(\cdot)$ is a non-linear activation function, $d_{i}$ is the degrees of unit $i$, $W_{2}$ is the learning weight matrix of GCN.
Then, we use $h_{i,4}$, $u^{i}_{i},u^{c}_{i}$ and $t_{i},z_{i}$ together to estimate $y_{i}$ of treated and control groups respectively, i.e.,
$\small\mu^{NN}(t_{i},z_{i},u^{i}_{i},u^{c}_{i},u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}})=\left\{\begin{aligned} MLP_{3}(z_{i},h_{i,2})&\quad t_{i}=1\\ MLP_{4}(z_{i},h_{i,2})&\quad t_{i}=0\end{aligned}\right.$
@@
$\mathcal{L}_{y}=\Sigma_{i=1}^{n}(y_{i}-\mu^{NN}(t_{i},z_{i},x_{i},x_{\mathcal{N}_{i}}))^{2}.$ (9)
Moreover, inspired by Jiang & Sun (2022); Cai et al. (2023), we further consider a balancing regularization term as our loss:
$\mathcal{L}_{IPM}=\text{IPM}(p(h_{i,4},t_{i},z_{i}),p(h_{i,4})p(t_{i})p(z_{i}))),$ (10)
where $\text{IPM}(p.q)=\sup_{g\in\mathcal{G}}|\int_{\mathcal{X}}g(x)(p(x)-q(x))dx|$ is the integral probability metric, measuring the distance between two distribution $p,q$, which can be implemented by Wasserstein Distance. Here the samples from $p(h_{i,4})p(t_{i})p(z_{i}))$ is obtained by randomly permuting $t_{i}$ and $z_{i}$ separately.
@@
$\mathcal{L}_{all}=-\textbf{ELBO}+\mathcal{L}_{sm}+\mathcal{L}_{y}+\mathcal{L}_{IPM}.$ (11)
6 Experiments

### v2

@@ IV Latent Confounder Identifiability Enables Networked Effect Identification
To begin with, following existing identifiable representation learning methods [39, 40], we first introduce the generative model as follows:
$p_{\bm{\theta}}(X,U|X_{\mathcal{N}})$ $=p_{\bm{f}}(X|U)p_{\bm{T},\bm{\lambda}}(U|X_{\mathcal{N}}),$ (2)
$p_{\bm{f}}(X|U)$ $=p_{\bm{\epsilon}}(X-\bm{f}(U)),$
@@
where $p_{\bm{\epsilon}}(X-\bm{f}(U))=p(\bm{\epsilon})$ is the distribution of additive noise $X=\bm{f}(U)+\bm{\epsilon}$. Further, we assume $p_{\bm{T},\bm{\lambda}}(U|X_{\mathcal{N}})$ follows the exponential family distribution as follows.
Assumption 6.
@@
where $\mathcal{Q}$ is the base measure, $\mathcal{C}$ is the normalizing constant. The $\bm{\lambda}(X_{\mathcal{N}})$ is an arbitrary function, and the sufficient statistics $\bm{T}(U)=[\bm{T}_{f}(U)^{T},\bm{T}_{MLP}(U)^{T}]^{T}$ contains a) the sufficient statistics $\bm{T}_{f}(U)^{T}=[\bm{T}_{1}(U_{(1)})^{T},\dots,\bm{T}_{1}(U_{(d_{U})})^{T}]$ of a factorized exponential family, where all the $\bm{T}_{i}(U_{(i)})$ have dimension larger or equal to 222 and $d_{U}$ is the dimension of $U$, and b) the output $\bm{T}_{MLP}(U)$ of a neural network with ReLU activations.
This assumption is introduced by [40]. The distribution in Assumption 6 is more flexible than the standard assumed distribution condition in identifiable representation learning (Eq. (7) in [39]). This assumption allows for the case that the different elements of latent confounders are not independent given the conditional set. The term $\bm{T}_{MLP}(U)$ does capture arbitrary dependencies between latent variables since the neural network with ReLU activation has universal approximation power.
@@
Suppose Assumption 6 holds, and suppose the following conditions hold: (1) The set $\{X\in\mathcal{O}|\varphi_{\bm{\epsilon}}(X)=0\}$ has measure zero where $\varphi_{\bm{\epsilon}}$ is the characteristic function of density $p_{\bm{\epsilon}}$. (2) $\bm{f}$ is injective and has all second-order cross derivatives. (3) The sufficient statistics in $\bm{T}_{\bm{f}}$ are all twice differentiable and each $\bm{T}_{\bm{f}}$ has at least one invertible dimension. (4) There exist $k+1$ distinct values $x_{\mathcal{N}_{0}},...,x_{\mathcal{N}_{k+1}}$ such that the matrix
$L=(\bm{\lambda}(x_{\mathcal{N}_{1}})-\bm{\lambda}(x_{\mathcal{N}_{0}}),.,\bm{\lambda}(x_{\mathcal{N}_{k+1}})-\bm{\lambda}(x_{\mathcal{N}_{0}}))$
@@
of size $k\times k$ is invertible where $k$ is the dimension of $\bm{T}$. Then we learn the true latent variable $U=\left[U^{i},U^{c},U^{n}\right]$ up to a permutation and element-wise transformations.
Proof is given in Appendix A.
@@
Discussion on assumptions: Assumption 6 indicates our theory holds for a rich family of conditional densities [41]. The assumption on the exponential family distribution is not strong, since many well-known distributions belong to this family, including the Gaussian, Uniform, Poisson distributions, and so on. Conditions (1)-(4) are mild and common assumptions in causal representation learning, e.g., [39, 40, 42]. In particular, Condition (3) is motivated by [42] to ensure that the identifiability holds for the recovered latent variables beyond their sufficient statistics. Among these, the most critical is Condition (4), which requires that the auxiliary information be sufficiently informative. In our networked setting, this translates to the existence of enough variability in neighbors’ covariates, a condition that is typically satisfied when a sufficiently rich set of covariates is collected, especially if some covariates are continuous.
Theorem 1 shows that, under mild assumptions, the latent confounders can be recovered up to an invertible transformation by utilizing the auxiliary information revealed by networked interference. Specifically, the recovered latent variables $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$ satisfy $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}=h(U^{i},U^{c},U^{n})$ for some bijective element-wise transformations $h$. However, the exact correspondence between the recovered dimensions and the original latent variables $U^{i},U^{c}$ and $U^{n}$ remains unidentified due to an unknown permutation. This ambiguity, nevertheless, can be resolved by exploiting the asymmetric roles that these latent variables play in the causal structure: $U^{i}$ only affects the unit itself, $U^{c}$ only affects the neighbors, and $U^{c}$ affects both, as shown in the following theorem.
Theorem 2.
Suppose Theorem 1 holds. Then, by maximizing the log-likelihood of $Y$, i.e., $\max\log p(Y\mid U^{i},U^{c},U^{c}_{\mathcal{N}},U^{n}_{\mathcal{N}}),$ the latent variables $U^{i}$, $U^{c}$, and $U^{n}$ are identifiable up to a permutation and element-wise transformations within each group, respectively.
Proof is given in Appendix B.
@@
Theorem 2 indicates that the three categories of recovered latent confounders can be further disentangled, i.e., the recovered $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$ satisfying $\hat{U}^{i}=h_{i}(U^{i}),\hat{U}^{c}=h_{c}(U^{c}),\hat{U}^{n}=h_{n}(U^{n})$ for some simple functions $h_{i}.h_{c},h_{n}$. Based on Theorem 2 above, we can further identify networked effect as follows:
Theorem 3.
Suppose Assumptions 1, 2, 3, 5, and Theorems 1 and 2 hold, the networked effects $\mu(t,z,x,x_{\mathcal{N}})$ can be identified by:
$\mu(t,z,x,x_{\mathcal{N}})$ (4)
$=\mathbb{E}[\mathbb{E}[Y\mid\hat{U}^{i}=\hat{u}^{i},\hat{U}^{c}=\hat{u}^{c},\hat{U}^{c}_{\mathcal{N}}=\hat{u}^{c}_{\mathcal{N}},\hat{U}^{n}_{\mathcal{N}}=\hat{u}^{n}_{\mathcal{N}},$
$\quad\quad\quad\quad\quad T=t,Z=z]]$
where $\hat{\circ}$ denotes the recovered latent variables, and similarly for the average effects $\psi(t,z)$.
Proof is given in Appendix C.
@@
Theorem 3 indicates that if we can identify latent confounders, the networked effect is thereby identifiable. This necessitates the utilization of identifiable representation learning techniques for causal inference in the presence of networked interference and latent confounders. Built upon the theoretical findings, in the next section, we propose our practical networked effect estimator that utilizes identifiable causal representation learning techniques.
@@
Figure 3: Model architecture of our proposed method named CaLaNet. The representation learning module aims to learn the latent confounders. The feature module aggregates the information of the confounders of unit $i$ and its neighbor. The outcome estimator module aims to estimate potential outcomes of unit $i$.
V Methodology
@@
In this section, we devise our networked effect estimator in the presence of latent confounders. Specifically, our estimator contains three modules, including the representation learning module, the feature module, and the outcome estimator module. The representation learning module is built on Theorem 1, aiming to correctly recover three types of latent confounders $u^{i},u^{c},u^{n}$. The feature module is built on Theorems 2 and 3, aggregating the information from units’ and neighbors’ information. This aggregated information is then input into the outcome estimator module to predict the networked causal effects. Overall, our model architecture is shown in Figure 3.
V-A Representation Learning Module
@@
Following existing work [43, 6, 25, 9], we use Graph Convolution Networks (GCN [44, 45]) to aggregate the information of covariates of unit $i$ and its neighbors, i.e., $x_{i},x_{\mathcal{N}_{i}}$:
$h^{neigh}_{i,1}=\sigma(\sum_{j\in\mathcal{N}_{i}}\frac{1}{\sqrt{d_{i}d_{j}}}W_{1}^{T}x_{j}),$ (5)
@@
where $\sigma(\cdot)$ is a nonlinear activation function, $d_{i}$ is the degrees of unit $i$, $W_{1}$ is the learning weight matrix of GCN, and $MLP_{1}$ is a multilayer perception (MLP). Here, the representation $h_{i,2}$ serves as $x_{i}$ and $x_{\mathcal{N}_{i}}$.
Then, given $x_{i}$ and $x_{\mathcal{N}_{i}}$, we employ the identifiable representation learning technique to recover latent confounders. Specifically, according to Assumption 6, we parametrize the prior following [40, 42]:
@@
where $u_{i}=[u_{i}^{i},u_{i}^{c},u_{i}^{n}]$ and $MLP_{2}(u_{i})$ serves as $\bm{T}_{MLP}$, the concatenated $[u_{i},u_{i}^{2}]$ serves as $\bm{T}_{f}$, $MLP_{3}(h_{i,2})$ serves as $\bm{\lambda}_{MLP}$, and $MLP_{4}(h_{i,2})$ serves as $\bm{\lambda}_{f}$ in Assumption 6.
@@
$q(u^{i}_{i},u^{c}_{i},u^{n}_{i}|x_{i},x_{\mathcal{N}_{i}})=\prod_{i=0}^{d_{U}}\mathcal{N}(\mu=\hat{\mu}_{u_{i}},\sigma^{2}=\hat{\sigma}_{u_{i}}^{2}),$ (7)
where $\hat{\mu}_{u_{i}}$ and $\hat{\sigma}_{u_{i}}^{2}$ are the mean and variance of the Gaussian distribution parametrized by MLPs using $h_{i,2}$ as input, and $d_{U}$ is the dimension of $U$.
@@
$p(x_{i}|u^{i}_{i},u^{c}_{i},u^{n}_{i})=\prod_{i=0}^{d_{X}}\mathcal{N}(\mu=\hat{\mu}_{x},\sigma^{2}=v_{x}^{2})$ (8)
or $p(x_{i}|u^{i}_{i},u^{c}_{i},u^{n}_{i})=\prod_{i=0}^{d_{X}}\mathbf{Bern}(\pi=\hat{\pi}_{x}),$
where for the continuous case $\hat{\mu}_{x}$ is the mean of the Gaussian distribution parametrized by an MLP using the sampled $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ from posterior as input and $v_{x}^{2}$ is the fixed variance of Gaussian distribution, and for the discrete case $\hat{\pi}_{x}$ is the mean of Bernoulli distribution similarly parametrized by an MLP, and here $d_{X}$ is the dimension of $X$.
In this module, we use the negative variational Evidence Lower BOund (ELBO) as the loss function, defined as
@@
Detailed ELBO devotions are given in Appendix D.
Since the prior belongs to a multivariate exponential family distribution with unknown normalization constant $\mathcal{C}$, it is infeasible to learn $\bm{T}$ and $\bm{\lambda}$ by optimizing the KL-divergence term in Eq. (9). Therefore, we utilize the widely used score matching technique [46] for training unnormalized densities to learn the parameters of $\bm{T}$ and $\bm{\lambda}$ by minimizing
@@
[…]
$\mathcal{L}_{rlm}={\mathbb{E}}_{q(u^{i},u^{c},u^{n}|x.x_{\mathcal{N}})}[-\log p(x|u^{i},u^{c},u^{n})+$ (11)
[…]
V-B Feature Module and Outcome Estimator Module
@@
where $\sigma(\cdot)$ is a nonlinear activation function, $d_{i}$ is the degrees of unit $i$, $W_{2}$ is the learning weight matrix of GCN. Here, $h_{i,3}^{neigh}$ serves as the neighbor latent confounders $[u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}]$ and $h_{i,4}$ serves as the whole confounders $[u^{i}_{i},u^{c}_{i},u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}]$, used to prediction the outcomes $y_{i}$.
Then, we use $h_{i,4}$ and $t_{i},z_{i}$ together to estimate $y_{i}$ of treated and control groups respectively, i.e.,
$\small\mu^{NN}(t_{i},z_{i},u^{i}_{i},u^{c}_{i},u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}})=\left\{\begin{aligned} MLP_{3}(z_{i},h_{i,4}),&\quad t_{i}=1\\ MLP_{4}(z_{i},h_{i,4}),&\quad t_{i}=0\end{aligned}\right.,$
@@
$\mathcal{L}_{y}=-\Sigma_{i=1}^{n}(y_{i}-\mu^{NN}(t_{i},z_{i},x_{i},x_{\mathcal{N}_{i}}))^{2}.$ (12)
Moreover, inspired by [25, 7], we further consider a balancing regularization term to learn a balanced representation as our loss:
$\mathcal{L}_{IPM}=\text{IPM}(p(h_{i,4},t_{i},z_{i}),p(h_{i,4})p(t_{i},z_{i})),$ (13)
where $\text{IPM}(p.q)=\sup_{g\in\mathcal{G}}|\int_{\mathcal{X}}g(x)(p(x)-q(x))dx|$ is the integral probability metric, measuring the distance between two distribution $p,q$, which can be implemented by Wasserstein Distance. Here the samples from $p(h_{i,4})p(t_{i},z_{i}))$ is obtained by randomly permuting the observed treatment pairs $t_{i}$ and $z_{i}$.
@@
$\mathcal{L}_{all}=\mathcal{L}_{rlm}+\alpha\mathcal{L}_{IPM}+\beta\mathcal{L}_{y},$ (14)
where $\alpha$ and $\beta$ are the hyperparameters that control the weights of the respective loss terms.
We denote our method as CaLaNet (Causal representation learning-based Latent-confounder recovery for Networked effect estimation).
VI Experiments